OpenStudy (anonymous):

a water balloon is tossed in the air with an upward velocity of 25ft/s. the height h(t)in ft after t seconds is given by the function h(t)=-16^2+25t+3. 1.,After how many seconds will the balloon hit the ground? (hint: Use the quadratic formula) 2.,What will the height be at t = 1 second? please help

OpenStudy (anonymous):

Hi What do you need help with today :)

OpenStudy (anonymous):

@trinajoy um this question please im having trouble find out the second question the first

OpenStudy (anonymous):

btw what grade it this

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

To find out how many seconds it would take until the balloon hits the ground, you would use the quadratic formula: −b±b2−4ac−−−−−−−√2a −162+25t+3=0−25±625+192−−−−−−−−√−32−25±817−−−√−32 x = 1.674 seconds (other solution is negative) Same thing for t = 1 second, except the equation is equal to 1 instead of 0: −162+25t+3=1−162+25t+2=0−25±625+128−−−−−−−−√−32−25±753−−−√−32 x = 1.638 (other solution is negative) 0 commentsAdd a Comment

OpenStudy (anonymous):

that's the answer to the second question (8th g)

OpenStudy (anonymous):

@london9321 you see the bottom one that is the answer to the second question I.638 is the answer to the second question

OpenStudy (anonymous):

thank you sooo much @trinajoy

OpenStudy (anonymous):

np btw are you new to study island @london9321

OpenStudy (anonymous):

if so \bbox[8pt, pink,border:5pt solid red ]{\color{black}{\LARGE\cal~\bigstar\heartsuit\dagger Welcome~to~OpenStudy!\bigstar\heartsuit\dagger{{\cal\color{#872af0}{\hspace-274.5pt{\bigstar\heartsuit\dagger Welcome~to~OpenStudy!\bigstar\heartsuit\dagger}}}}~~}}

OpenStudy (anonymous):

if you are new to open study here

OpenStudy (anonymous):

\bbox[8pt, pink,border:5pt solid red ]{\color{black}{\LARGE\cal~\bigstar\heartsuit\dagger Welcome~to~OpenStudy!\bigstar\heartsuit\dagger{{\cal\color{#872af0}{\hspace-274.5pt{\bigstar\heartsuit\dagger Welcome~to~OpenStudy!\bigstar\heartsuit\dagger}}}}~~}}

OpenStudy (anonymous):

yea im knew ??